# Cryogenic source of atomic tritium for precision spectroscopy and neutrino-mass measurements

### v1

@@ III.2 Deuterium and tritium lfs
Technically, cooling the dissociator to 0.2 K can be done by thermally anchoring its chamber to the cold plate of the dilution refrigerator, typically operating at 0.1-0.2 K. Working with pure 4He requires higher temperatures K, which can be obtained by using a 3He refrigerator and cooling the next stages of the transfer line with a dilution refrigerator. In the Turku lab, a cryogenic system with two refrigerators in one cryostat: dilution and 3He type, was used for H hfs studies. The 3He refrigerator also has substantially larger cooling power than the dilution refrigerator cold plate.
@@
One strategy to increase the atom flux from the source is to increase the discharge RF power. As a reference, the H source in Turku is running at an average power of 1 mW as reported previously Semakin et al. (2025). The flux of atoms saturated at around this power value, and increasing it further by a factor of 2 stopped the flux completely. As a possible explanation, we consider an insufficient backflow of superfluid film into the dissociator chamber that could not be compensated for by evaporation of helium inside, leading to the dissociator drying out of the helium film at RF powers exceeding some critical value of mW. This problem can be solved by increasing the diameter of the transfer line and by coating the surface of the inner walls with a sintered layer of copper or silver powder. Then, the maximum dissipated power in the dissociator will be limited by the available cooling power at 0.2 or 0.4 K. In the latter case, using a 3He refrigerator, it can be mW, which may potentially increase the atomic flux by a factor of 50 and reach values well above s-1.
@@
Using 3He-4He films may be challenging when running the dissociator at the highest power of 50 mW. It is known that the transport properties of such films and the critical superflow velocity are reduced due to the presence of 3He, which contributes to a normal component of the film Hallock (1995). In this case, the concentration of 3He can be reduced to a level required for stable operation of the dissociator, or one may fall back to the operation with pure 4He films.
@@
Self-dissociation of T2 in the dissociator chamber by -decay electrons. So far, we have considered that the dissociation of molecules in the solid T2 layer in the dissociator chamber is performed by the electrons of the RF discharge. As described in section II.5, electrons resulting from the -decay of T also effectively produce atoms in the solid layer. In the previous work of the Turku group with T in T2 Sheludiakov et al. (2017), very thin films of a maximum of 250 nm were used, limited by the exemption limit for work with radioactive materials (1 GBq). We found that each -electron produced about 50 atoms. With a dissociation energy of 4.6 eV, this corresponds to a production efficiency of The penetration depth of 5.7 keV electrons in solid hydrogen is about m Schou and Sørensen (1978), and we expect that the dissociation number per T decay will be substantially larger for thicker films.
@@
We consider the following design for large flux applications. The cylinder-shaped dissociator chamber with a diameter of 10 cm and a length of 10 cm will be covered inside by a T2 layer of 1 m thickness. The total amount of T2 is moles and the rate of -decay events in this layer is s-1. Taking the same dissociation efficiency of 50 atoms per T decay, we evaluate the lower bound for the production rate of atoms in sold T2 as Ṅ s-1. If we assume that the efficiency scales linearly with thickness, then it should increase to 200 events/decay, and the T production rate will be s-1. This is larger than the T flux we may get from the dissociator by running the pulsed RF discharge above the solid film with 50 mW power. The heat released by the amount of T in the estimate above is evaluated as mW, five times smaller than the RF power. The self-dissociation of T2 seems to be a very efficient mechanism and will provide better efficiency than the RF discharge. As we mentioned in Section II E, the presence of the superfluid helium film improves the cooling of the T2 films and helps to avoid thermal explosions and partial evaporation of the film. One may try to increase the thickness of T2 further if the cooling by the helium film will still be able to stop thermal explosions. Finally, both dissociation methods, running the RF discharge on top of the T2 film, and self-dissociation, will work together, and reaching the flux of atoms above s-1 seems quite realistic.
@@ IV Conclusions
We proposed a concept of a cryogenic source of atomic tritium based on a pulsed RF discharge below 1 K, a technique successfully used to obtain large fluxes of atomic hydrogen. We expect an enhanced dissociation rate due to contributions of electrons resulting from the -decay of T. Working with T requires to prevent the interaction of atoms with the walls where they adsorb and recombine, even if a superfluid helium film is used. We suggest using a buffer gas cooling technique to provide a thermal link between atoms and the walls without physical contact. The atoms are expelled from the walls by magnetic field gradients. Thermal contact via the buffer gas may be easily adjusted by small changes in the temperature distribution in the gas transfer line. Optimal geometry and operating parameters may be found in experimental tests of the source prototype. We evaluate that the fluxes of T atoms exceeding s-1 can be reached at temperatures of mK, optimal for magnetic trapping and other experiments with the beams of slow D and T atoms.

### v2

@@ III.2 Deuterium and tritium lfs
Technically, cooling the dissociator to 0.2 K can be done by thermally anchoring its chamber to the cold plate of the dilution refrigerator, typically operating at 0.1-0.2 K. Working with pure 4He requires higher temperatures K, which can be obtained by using a 3He refrigerator and cooling the next stages of the transfer line with a dilution refrigerator. In the Turku lab, a cryogenic system with two refrigerators in one cryostat: dilution and 3He type, was used for H hfs studies. The 3He refrigerator also has substantially larger cooling power at 0.3 K than the dilution refrigerator cold plate.
@@
One strategy to increase the atom flux from the source is to increase the discharge RF power. As a reference, the H source in Turku is running at an average power of 1 mW as reported previously [25]. The flux of atoms saturated at around this power value, and increasing it further by a factor of 2 stopped the flux completely. As a possible explanation, we consider an insufficient backflow of superfluid film into the dissociator chamber that could not be compensated for by evaporation of helium inside, leading to the dissociator drying out of the helium film at RF powers exceeding some critical value of mW. This problem can be solved by increasing the diameter of the transfer line and by coating the surface of the inner walls with a sintered layer of copper or silver powder. Then, the maximum dissipated power in the dissociator will be limited by the available cooling power at 0.2 or 0.4 K. In the latter case, using a 3He refrigerator, it can be mW, which may potentially increase the atomic flux by a factor of 50 and reach values well above s-1.
@@
Using 3He-4He films may be challenging when running the dissociator at the highest power of 50 mW. It is known that the transport properties of such films and the critical superflow velocity are reduced due to the presence of 3He, which contributes to a normal component of the film [78]. In this case, the concentration of 3He can be reduced to a level required for stable operation of the dissociator; the operation may resume with pure 4He films.
III.3 Self-dissociation of T2 in the dissociator chamber by -decay electrons.
@@
So far, we have considered that the dissociation of molecules in the solid T2 layer in the dissociator chamber is performed by the electrons of the RF discharge. As described in Section II.5, electrons resulting from the -decay of T also effectively produce atoms in the solid layer. In the previous work of the Turku group with T in T2 [60], very thin films of a maximum of 250 nm were used, limited by the exemption limit for work with radioactive materials (1 GBq). We found that each -electron dissociated about 25 molecules and produced 50 atoms. With a dissociation energy of 4.6 eV, this corresponds to a production efficiency of The penetration depth of 5.7 keV electrons in solid hydrogen is about m [79], and we expect that the dissociation number per T decay will be substantially larger for thicker films.
@@
We consider the following design for large flux applications. The cylinder-shaped dissociator chamber with a diameter of 10 cm and a length of 10 cm will be covered inside by a T2 layer of 1 m thickness. The total amount of T2 is moles and the rate of -decay events in this layer is s-1. Taking the same dissociation efficiency of 50 atoms per T decay, we evaluate the lower bound for the production rate of atoms in solid T2 as Ṅ s-1. If we assume that the efficiency scales linearly with thickness, then it should increase to 200 events/decay, and the T production rate will be s-1. This is larger than the T flux we may get from the dissociator by running the pulsed RF discharge above the solid film with 50 mW power. The heat released by the amount of T in the estimate above is evaluated as mW, five times smaller than the RF power. The self-dissociation of T2 seems to be a very efficient mechanism and may provide better efficiency than the RF discharge. As we mentioned in Section II E, the presence of the superfluid helium film improves the cooling of the T2 films and helps to avoid thermal explosions and partial evaporation of the film. One may try to increase the thickness of T2 further if the cooling by the helium film will still be able to stop thermal explosions. Finally, both dissociation methods, running the RF discharge on top of the T2 film, and self-dissociation, will work together, and reaching the flux of atoms above s-1 seems quite realistic.
@@ IV Conclusions
We propose a concept of a cryogenic source of atomic tritium based on a pulsed RF discharge below 1 K, a technique successfully used to produce large fluxes of atomic hydrogen. We expect an extra enhancement of the dissociation rate due to contributions of electrons resulting from the -decay of T. Working with T requires to prevent the interaction of atoms with the walls where they adsorb and recombine, even if a superfluid helium film is used. We suggest using a buffer gas cooling technique to provide a thermal link between atoms and the walls without physical contact. The atoms are expelled from the walls by magnetic field gradients. Thermal contact via the buffer gas may be easily adjusted by small changes in the temperature distribution in the gas transfer line. Optimal geometry and operating parameters may be found in experimental tests of a source prototype. We evaluate that the fluxes of T atoms exceeding s-1 can be reached at temperatures of mK, optimal for magnetic trapping and other experiments with the beams of slow D and T atoms.
V Acknowledgments
This project was supported by the Jenny and Antti Wihuri foundation.